Observation of Top Quark

Production

Pekka K. Sinervo University of Toronto / IPP

1. Review of Prior Results

2. Description of new CDF experiment

3. Results of CDF Search

4. Outlook & Conclusions

University of Toronto March 2, 1995

Introduction: Why a Top Quark?

The Standard Model of Elementary Particle Physics

• Three “generations” of quarks and leptons

( \begin{matrix}u \\ d \end{matrix} ) ( \begin{matrix}c \\ s \end{matrix} ) ( \begin{matrix}t \\ b \end{matrix} )
( \begin{matrix}e \\ ν_{e} \end{matrix} ) ( \begin{matrix}μ \\ ν_{μ} \end{matrix} ) ( \begin{matrix}τ \\ ν_{τ} \end{matrix} )

“Isospin” doublets that dictate particle interactions

• Forces carried by bosons

Electromagnetic force  γ
Strong force  eight gluons
Weak force  W ±, Z∘ bosons
Gravity  gravitons

SM Predictions for Top Quarks

The asymmetry in e^{+} e^{−} → bb̄ scattering implies that b is part of SU(2) doublet

• top quark has standard SU(2) couplings

• CKM matrix predicts dominant decay t → W + b

Corrections to W^{±} and Z^{∘} masses depend on heaviest fermion

• LEP, SLD measurements on the Z resonance require

Mtop = 178 ± 10{+ 18 \atop −19} GeV/c2

D. Schaile, CERN-PPE/94-162

CDF Reported Evidence for Top in 1994

CDF reported observation of 2 dilepton events

• Expected ∼ 0.5 events in this channel

Observed 10 lepton+jet events with b quark “tag”

• Expected 5.4 events background

• 7 events fully reconstructed

D0 / had observed excess of events

• found 18 events when expected ∼ 8 from background

What Does it Take to Discover Top?

• Observe “anomalous” production of W^{+} W^{−} events

– Use dilepton events – both W’s decay leptonically

• Observe W^{+} W^{−} events associated with a b quark

– Use lepton + jet events

• Show events come from heavy fermion decays

• Confirm rate is consistent with t t production

– Measure cross section using tagged events

1992-95 Collider Run

CDF and D0 / have collected ×3 more data

• Run 1A (1992-93) provided ∼ 20 pb−1

• Run 1B (1994-95) has produced another ∼ 55 pb−1 – D0 / had updated results on top search (+22 pb−1)

• Will report on ∼ 65 pb−1 of CDF data

By far the largest and most sensitive dataset to top

• For a top quark with a mass of 160 GeV/c2 – expect to observe 3-4 dilepton events

– expect to observe 30 lepton + jet events

Dilepton Top Quark Search

t t decaying into 2 leptons create unique signature

• Two charged leptons from W ± decay

• Imbalance in total energy due to undetected neutrinos • Two jets from bottom quarks

CDF performed a search requiring

• Two clean, high PT lepton candidates

• Missing transverse energy E̸T > 25 GeV

• Require two jets with ET > 10 GeV and |η| < 2.0 • Reject specific backgrounds (e.g. Z∘)

Dielectron Data for Run 1A and 1B

8

DiMuon Data for Run 1A and 1B

9

Electron-Muon Data for Run 1A and 1B

10

Summary of Dilepton Search

Observe a total of 6 candidate events

• Observe 1 μ+ μ− event and 5 eμ events

• Expect 1.3 ± 0.3 events from background

Background  Expected Background Rate
Z → τ+ τ−  0.38 ± 0.07
One or both leptons fake  0.23 ± 0.15
bb̄  0.03 ± 0.02
W W, W Z and ZZ  0.21 ± 0.07
Drell-Yan  0.44 ± 0.28
Total  1.3 ± 0.3

Probability of background fluctuation is 3.3 × 10^{−3}

Lepton + Jets Search Philosophy

Decay mode with the greatest sensitivity is “lepton+jets”

• About 30% of t t events decay this way

• Look for W → lνl decay

• Require two jets from W → qq̄ decay

• Require at least one more jet from b quarks – must “tag” at least one b quark jet

Selection of W candidates straightforward

• Require high PT isolated lepton candidate

• Require significant E̸T

• Require three or more high PT jets

Have now quadrupled our sample of W + ≥ 3 jet events

• Have 203 events before b tagging

• Expect ∼ 30 to come from t t production

B Tagging Techniques

Use two properties of b quarks to identify them

1. About 20% of b’s decay via

b → e−νeX and b → μ−νμX

• Look for jets with “soft” leptons in them

• Call these soft-lepton tags (SLT’s)

• Expect 15% of t t events to have at least one SLT tag 2. “Long” b lifetime – cτ ∼ 450 microns

• CDF uses Silicon Vertex Detector to identify jets with a secondary vertex

• We call these “SVX” tags

• Expect 42% of t t events to have at least one SVX tag

Idea is to just “count” number of tagged events

b Quark Tagging in W+Jet Events

Applied these tagging techniques to W+jet events

• The W + 1 and W + 2 jet samples are control samples – Don’t expect significant top contribution

– Theory predicts low b hadron content

Sample  SVX bkg SVX tags  SLT bkg  SLT tags
W+1 jet  50 ± 12  40  159 ± 25  163
W+2 jet  21 ± 7  34  46 ± 7  55
W+≥ 3 jet  6.7 ± 2.1  27  15.4 ± 2.3  23

• Probability of background fluctuation is 1 × 10−6 – Combining SVX, SLT and DIL channels

• Agrees well with expected rate from t t production

SVX Tagging Results

The SVX tagged sample is single most significant observation

• Observe 27 tags when we expect 6.7 ± 2.1 background • Probability of background fluctuation 2 × 10−5

Measurement of Top Mass

We use event kinematics to estimate top quark mass

• Require events to have a fourth jet – Require ET > 8 and |η| < 2.4 for 4rth jet – High efficiency (∼ 85% for t t)

• Combine all possible jet pairs to form W → qq̄ – Choose one with best χ2

– Require χ2 < 10

• Use b-tag information to identify b quark jet

• Fit events requiring kinematic constraints

This selection leaves us with 19 tagged events

• Estimate background to be 6.9{+2.5 \atop −1.9} events – Primarily QCD W+jet

Procedure is dependent on theory predictions

Result of Mass Fits

The fitted mass for the 19 events is consistent with t t

• Probability that background would fluctuate to give peak with similar significance is < 2 × 10−2

M_{top} = 176 ± 8 ± 10 GeV/c^{2}

Physics Comments

Most interesting immediate result is how heavy top appears to be

• In good agreement with indirect measurements

• Consistent with MW measurements and SM

• Cross section is somewhat high

Cross Section Measurement

Use the channel with least background to estimate σ

• Subtract backgrounds from SVX-tags – Estimated background 5.5 ± 1.8 events

• Correct for tagging efficiency and acceptance – Efficiency is 0.0338 ± 0.0087

σ_{tt̄} = 6.8_{−2.4}^{+3.6} pb

Consistency Checks Revisited

Two checks last year raised concern

• Rate of W + 3 jet versus W + 4 jet events – Didn’t see “enough” W + 4 jet events

• Saw two Z + 4 jet b-tagged events – expected 0.6

Both appear to have been statistical fluctuations

1. Predicted rate of W+4 jet events now in good agreement

Process  Observed Events
Predicted QCD W+4 jet rate  13 ± 6.5
Observed W+4 jet rate  25 ± 6

2. Have not observed more tagged Z + 4 jet events

Number of Jets  Observed Tags  Predicted
1 jet  15  17.5
2 jet  3  4.2
≥ 3 jets  2  1.5

Prospects for Run 1A + Run 1B

Expect statistical uncertainty on M_{top} to scale as 1/√N

• Have performed MC “toy” studies to evaluate this

• Expect statistical σM ∼ 6 GeV/c2

• Systematic σM hardest to control

– Calibrate gluon radiation using other processes

∗ Study of γ+ jet events allow us to calibrate energy

flow out of jets

– Sensitivity to background shape also should go as 1/√N ∗ Will have enough statistics to measure shape of non-

W background

• Reasonable to expect systematic uncertainty to drop as 1/√N

Can begin to estimate couplings

• Will have of order 10 dilepton events

• Expect of order 50 lepton + jet events

• Might expect ∼ 30% measurement of \frac{BR(top → leptons )}{BR(top → hadrons )}

Conclusions

CDF has conclusively observed top quark production

• All three channels show significant signals

• Kinematic analysis confirms interpretation

• Probability of background fluctuation: – Counting experiment significance 1 × 10−6 – Mass analysis significance < 2 × 10−2

– Combined significance is 3.7 × 10−7

Best estimate of top quark mass is M_{top} = 176 ± 8 ± 10 GeV/c^{2}

Result is confirmed by D0 /

• Have no official details as yet

Continuing data collection for more detailed top studies